Bombs Away Video Games Corporation has forecasted the following monthly sales January $107,000 July February100,000 August March $ 52,000 52,000 32,000 September 62,000 92,000 27,000 November 112,000 42,000 December 130,000 ril May June 32,000 October Total annual sales $840,000 Bombs Away Video Games sells the popular Strafe and Capture video game. It sells for $5 per unit and costs $2 per unit to produce. A level production policy is followed. Each months production is equal to annual sales (in units) divided by 12. Of each months sales, 40 percent are for cash and 60 percent are on account. All accounts receivable are collected in the month after the sale is made a. Construct a monthly production and inventory schedule in units. Beginning inventory in January is 32,000 units Bombs Away Video Games Corporation Production and Inventory Schedule in Units Beginning+Production-|Sales Invento Invento Ending Janua February March April Ma June 32,000 August September October November December b. Prepare a monthly schedule of cash receipts. Sales in December before the planning year are $100,000 Bombs Away Video Games Corporation Cash Receipts Schedule Februa Janua March April Ma June Sales Cash receipts Cash sales Prior months credit sales Total cash receipts $Bombs Away Video Games Corporation Cash Receipts Schedule August September October November December Sales Cash receipts Cash sales Prior months credit sales Total cash receipts 0 0 0 c. Prepare a cash payments schedule for January through December. The production costs of $2 per unit are paid for in the month in which they occur. Other cash payments, besides those for production costs, are $52,000 per month. Bombs Away Video Games Corporation Cash Payments Schedule Janua Februa March Ma June Production cost Other cash Total cash ents 0 0 Bombs Away Video Games Corporation Cash Payments Schedule Jul August September October November December Production cost Other cash Total cash payments d. Prepare a monthly cash budget for January through December using the cash receipts schedule from part b and the cash payments schedule from part c. The beginning cash balance is $5,000, which is also the minimum desired. (Negative amounts should be indicated by a minus sign.) Bombs Away Video Games Corporation Cash Budget January Februa March April Ma June Beginning cash Net cash flow Cumulative cash balance Monthly loan or re Ending cash balance Cumulative loan balance Bombs Away Video Games Corporation Cash Budget August September October November December inning cash Net cash flow Cumulative cash balance Monthly loan or re Ending cash balance Cumulative loan balance
Answer #1
Particulars January February March April May June July August September October November December
Sales Revenue 1,07,000.00 1,00,000.00 32,000.00 32,000.00 27,000.00 42,000.00 52,000.00 52,000.00    62,000.00 92,000.00 1,12,000.00 1,30,000.00
Sales Unit @5 PU 21400 20000 6400 6400 5400 8400 10400 10400 12400 18400 22400 26000
Production PM (Total Sales/12) 14000 14000 14000 14000 14000 14000 14000 14000 14000 14000 14000 14000
Production Cost @ 2 PU      28,000.00      28,000.00 28,000.00 28,000.00 28,000.00 28,000.00 28,000.00 28,000.00    28,000.00 28,000.00      28,000.00      28,000.00

Solution A

Bombs Away Video Games Corporation
Production & Inventory Schedules in Units
Beginning Inventory + Production - Sales = Ending Inventory
January 32000 + 14000 - 21400 = 24,600.00
February 24600 + 14000 - 20000 = 18,600.00
March 18600 + 14000 - 6400 = 26,200.00
April 26200 + 14000 - 6400 = 33,800.00
May 33800 + 14000 - 5400 = 42,400.00
June 42400 + 14000 - 8400 = 48,000.00
July 48000 + 14000 - 10400 = 51,600.00
August 51600 + 14000 - 10400 = 55,200.00
September 55200 + 14000 - 12400 = 56,800.00
October 56800 + 14000 - 18400 = 52,400.00
November 52400 + 14000 - 22400 = 44,000.00
December 44000 + 14000 - 26000 = 32,000.00


Solution B

Bombs Away Video Games Corporation
Cash Receipts Schedule
January February March April May June July August September October November December
Sales 1,07,000.00 1,00,000.00 32,000.00 32,000.00 27,000.00 42,000.00 52,000.00 52,000.00    62,000.00 92,000.00 1,12,000.00 1,30,000.00
Cash Receipts
Cash Sales @ 40%      42,800.00      40,000.00 12,800.00 12,800.00 10,800.00 16,800.00 20,800.00 20,800.00    24,800.00 36,800.00      44,800.00      52,000.00
Prior Month's Credit Sales 60000      64,200.00 60,000.00 19,200.00 19,200.00 16,200.00 25,200.00 31,200.00    31,200.00 37,200.00      55,200.00      67,200.00
Total Receipts 102800 104200 72800 32000 30000 33000 46000 52000 56000 74000 100000 119200

Solution C

Bombs Away Video Games Corporation
Cash Payment Schedule
January February March April May June July August September October November December
Production Cost @ 2 PU 28000 28000 28000 28000 28000 28000 28000 28000 28000 28000 28000 28000
Other Cash Payment 52000 52000 52000 52000 52000 52000 52000 52000 52000 52000 52000 52000
Total Cash Payment 80000 80000 80000 80000 80000 80000 80000 80000 80000 80000 80000 80000

Solution D

Bombs Away Video Games Corporation
Cash Budget
January February March April May June July August September October November December
Beginning Cash 5000 27800 52000 44800 5000 5000 5000 5000 5000 5000 5000 5000
Net Cash Flow 22800 24200 -7200 -48000 -50000 -47000 -34000 -28000 -24000 -6000 20000 39200
Cumulative Cash Balance 27800 52000 44800 -3200 -45000 -42000 -29000 -23000 -19000 -1000 25000 44200
Monthly Loan or (Repayment) 0 0 0 -8200 -50000 -47000 -34000 -28000 -24000 -6000 20000 39200
Ending Cash Flow 27800 52000 44800 5000 5000 5000 5000 5000 5000 5000 5000 5000
Cumulative Loan Balance 0 0 0 -8200 -58200 -105200 -139200 -167200 -191200 -197200 -177200 -138000